3. It is estimated that 2% of the meinbers of a certain population are infected with Hepatitis C virus. A diagnostic test for detecting Hepatitis C yields 4% false positive and 5% false negative results. (a) Find the probability that a person randomly chosen from this popu- lation has a positive result of the diagnostic test for Hepatitis C (b) Find the probability that a random person who has a positive result of the diagnostic test, actually has Hepatitis C.
Answer #1

a)

P(tested positive)=P(has Hepatitis C and tested positive)+P(not has Hepatitis C and tested positive)

=0.02*(1-0.05)+(1-0.02)*0.04=0.0582

b)

P(has Hepatitis C given tested positive)=P(has Hepatitis C and tested positive)/P(tested positive)

=0.02*(1-0.05)/0.0582=0.32646